# MdEval: Massively Multilingual Code Debugging

### v1

@@ 3.1 Experiment Setup
DS-Coder-Instruct 6.7B 48.3 45.3 47.1 46.7 50.8 48.5 48.2 50.0 45.0
CodeQwen1.5-Chat 7B 65.3 68.8 75.7 51.7 63.1 68.2 77.1 84.3 68.3
Qwen2.5-Coder-Instruct 7B 80.0 67.2 90.0 68.3 83.1 74.2 73.5 87.1 85.0
Yi-Coder-Chat 9B 67.1 59.4 70.0 58.3 67.7 63.6 56.6 80.0 70.0
DS-Coder-V2-Lite-Instruct 16B 63.7 70.3 65.7 56.7 60.0 66.7 56.6 71.4 61.7
xDebugGen (Our Method) 7B 60.4 59.4 61.4 68.3 56.1 67.7 65.1 65.7 61.7
Method Json Julia PHP Pascal Python R Ruby Rust Scala Swift
Closed-Source Models
GPT-4o-240806 91.5 91.0 92.2 93.8 98.4 98.3 90.5 94.3 95.3 95.0
GPT-4o-mini-240718 78.0 92.5 89.1 95.4 93.4 95.0 90.5 90.0 82.8 86.7
Claude-3.5-sonnet-240620 89.8 94.0 95.3 95.4 95.1 95.0 95.2 97.1 90.6 96.7
Claude-3.5-sonnet-241022 89.8 92.5 95.3 95.4 98.4 98.3 93.7 100.0 95.3 98.3
1B+ Models
DS-Coder-Instrcut 50.8 52.2 51.6 56.9 49.2 56.7 63.5 55.7 42.2 55.0
Yi-Coder-Chat 62.7 58.2 56.2 70.8 73.8 53.3 41.3 71.4 42.2 50.0
Qwen2.5-Coder-Instruct 52.5 56.7 50.0 47.7 44.3 46.7 57.1 52.9 51.6 53.3
6B+ Models
DS-Coder-Instruct 42.4 44.8 51.6 50.8 52.5 55.0 42.9 50.0 45.3 51.7
CodeQwen1.5-Chat 59.3 53.7 60.9 43.1 72.1 65.0 77.8 62.9 60.9 63.3
Qwen2.5-Coder-Instruct 61.0 85.1 95.3 76.9 73.8 88.3 81.0 74.3 79.7 96.7
Yi-Coder-Chat 52.5 74.6 84.4 63.1 68.9 53.3 69.8 75.7 64.1 75.0
DS-Coder-V2-Lite-Instruct 54.2 61.2 62.5 76.9 62.3 55.0 69.8 68.6 57.8 68.3
xDebugGen (Our Method) 55.9 62.7 60.9 52.3 52.5 60.0 60.3 55.7 54.7 66.7
@@
Model Size Avgall C C# CLISP CPP F# Go JS Java
@@
DS-Coder-Instrcut 1.3B 18.4 14.1 20.0 20.0 18.2 9.2 8.4 30.0 23.3
Yi-Coder-Chat 1.5B 1.4 3.1 2.9 3.3 3.0 0.0 0.0 0.0 0.0
Qwen2.5-Coder-Instruct 1.5B 1.2 3.1 2.9 3.3 3.0 1.5 0.0 0.0 0.0
6B+ Models
DS-Coder-Instruct 6.7B 9.3 10.9 2.9 6.7 12.1 0.0 6.0 28.6 8.3
CodeQwen1.5-Chat 7B 5.7 4.7 11.4 10.0 4.5 1.5 1.2 8.6 8.3
Qwen2.5-Coder-Instruct 7B 8.5 12.5 8.6 6.7 12.1 0.0 12.0 11.4 6.7
Yi-Coder-Chat 9B 14.5 25.0 14.3 11.7 18.2 3.1 9.6 18.6 18.3
DS-Coder-V2-Lite-Instruct 16B 13.6 14.1 14.3 8.3 7.6 18.5 14.5 11.4 5.0
xDebugGen (Our Method) 7B 4.6 7.8 8.6 6.7 7.6 0.0 2.4 4.3 1.7
Method Json Julia PHP Pascal Python R Ruby Rust Scala Swift
@@
DS-Coder-Instrcut 10.2 20.9 31.2 9.2 34.4 15.0 33.3 7.1 12.5 15.0
Yi-Coder-Chat 5.1 0.0 0.0 1.5 3.3 0.0 0.0 2.9 0.0 0.0
Qwen2.5-Coder-Instruct 1.7 0.0 0.0 0.0 3.3 0.0 0.0 2.9 0.0 0.0
6B+ Models
DS-Coder-Instruct 6.8 7.5 4.7 1.5 19.7 5.0 25.4 11.4 7.8 1.7
CodeQwen1.5-Chat 6.8 1.5 0.0 6.2 9.8 1.7 14.3 10.0 1.6 0.0
Qwen2.5-Coder-Instruct 15.3 3.0 1.6 6.2 24.6 6.7 12.7 7.1 1.6 3.3
Yi-Coder-Chat 6.8 19.4 6.2 6.2 36.1 8.3 17.5 15.7 14.1 11.7
DS-Coder-V2-Lite-Instruct 3.4 17.9 12.5 12.3 16.4 25.0 15.9 12.9 15.6 20.0
xDebugGen (Our Method) 5.1 3 0.0 1.5 14.8 3.3 9.5 5.7 0.0 0.0
@@ 3.2 Main Results
Table 3 presents the Pass@1 results of different models on MdEval for the multilingual automated program repair task (given question with buggy code, request the model to fix buggy code). The results indicate a marked disparity between closed-source state-of-the-art models and the majority of open-source models across nearly all programming languages. Notably, GPT-4o, Claude-3.5-sonnet, and DS-Coder-V2-236B excel in this task and demonstrate significant performance advantages over other models. Furthermore, our baseline model xDebugCoder, was fine-tuned using only 16K bug-related data MdEval-Instruct. Despite the limited size of this dataset, the model demonstrated competitive performance compared to others of similar scale, highlighting the effectiveness of MdEval-Instruct in enhancing the debugging capabilities of models.
Code Review.
Table 4 displays the accuracy for multilingual code review tasks. The results show that closed-source models still significantly outperform open-source models in the code review task. This task requires the model to distinguish between buggy and correct code. The closed-source models demonstrate a strong ability to understand complex code logic, achieving an accuracy rate of approximately 90%. In contrast, the smaller open-source model exhibits significant challenges, with an accuracy rate of around 50%. This disparity underscores the limitations of the current open-source options in effectively interpreting intricate coding patterns.
@@
In the bug identification task, the goal is to identify the error type in a given code snippet, where the LLMs analyze source code for defects and choose the correct bug type from the pre-defined 39 bug types. Table 5 lists the all results of the bug identification. Notably, the closed-source LLMs, such as GPT-4o and Claude series, have the dominant advantages, outperforming the open-source LLMs by nearly +10 points. Bug identification with 39 bug types poses a daunting challenge to the LLMs, requiring alignment capability of LLMs between the given code snippet and its corresponding bug type.
@@ 4 Further Analysis
As shown in Figure 6, The performance of models on the automated program repair task varies across different error types, highlighting the strengths and weaknesses of these models in addressing specific challenges. Consistently, the models demonstrate robust capabilities in repairing syntax errors, reference errors, and logic errors. These error types tend to be more straightforward and well-defined, allowing the models to leverage their knowledge effectively to identify and correct issues with high accuracy. In contrast, the models exhibit their worst performance when dealing with language-specific errors. Language-specific errors can arise from unique syntax rules, idiomatic expressions, or even cultural programming practices that are not universally applicable. As a result, addressing these types of errors presents a significant challenge and underscores the need for further improvements in the models’ training processes and methodologies.

### v2

@@ 3.1 Experiment Setup
DS-Coder-Instruct 6.7B 56.3 37.1 60.9 56.7 63.6 60.0 56.6 8.6 61.3 75.0 23.7 64.2 6.9 52.2 60.0 78.7 51.7 88.9 80.0 60.9 68.3
CodeQwen1.5-chat 7B 42.6 34.3 34.4 43.3 33.3 41.5 42.2 10.3 54.7 55.0 20.3 62.7 8.6 49.3 41.5 52.5 30.0 69.8 62.9 34.4 61.7
CodeLlama-Instruct 7B 27.2 2.9 20.3 25.0 25.8 24.6 22.9 19.0 53.3 6.7 15.3 37.3 12.1 24.6 33.8 42.6 16.7 50.8 48.6 14.1 41.7
CodeGemma-Instruct 7B 45.9 34.3 32.8 3.3 43.9 44.6 44.6 19.0 60.0 68.3 25.4 64.2 0.0 56.5 36.9 65.6 40.0 73.0 67.1 56.2 70.0
Qwen2.5-Instruct 7B 50.4 57.1 47.8 38.3 50.7 61.5 26.5 8.6 60.0 81.7 32.2 61.2 8.3 73.9 47.7 70.5 50.0 58.7 62.9 60.9 45.0
Qwen2.5-Coder-Instruct 7B 61.7 58.6 60.9 61.7 60.6 70.8 47.0 19.0 60.0 81.7 37.3 74.6 22.4 73.9 61.5 77.0 65.0 73.0 78.6 70.3 75.0
OpenCoder-Instruct 8B 53.4 10.0 56.2 46.7 50.0 66.2 8.4 13.8 66.7 78.3 27.1 74.6 15.5 62.3 53.8 77.0 61.7 76.2 81.4 71.9 75.0
Meta-Llama-3-Instruct 8B 37.9 51.4 35.8 8.3 42.0 30.8 21.7 10.3 60.0 41.7 20.3 49.3 0.0 55.1 40.0 57.4 50.0 49.2 48.6 46.9 28.3
Meta-Llama-3.1-Instruct 8B 42.1 57.1 41.8 26.7 40.6 44.6 21.7 6.9 56.0 60.0 20.3 49.3 8.3 65.2 32.3 50.8 40.0 58.7 61.4 53.1 38.3
Yi-Coder-Chat 9B 50.6 45.7 54.7 28.3 47.0 40.0 42.2 22.4 65.3 76.7 20.3 58.2 3.4 52.2 58.5 65.6 45.0 68.3 68.6 71.9 68.3
14B+ Models
Qwen2.5-Instruct 14B 57.7 58.6 62.7 61.7 66.7 60.0 21.7 13.8 62.7 78.3 28.8 59.7 10.0 69.6 66.2 80.3 68.3 74.6 77.1 76.6 56.7
DS-Coder-V2-Lite-Instruct 2.4/16B 56.7 10.0 56.2 43.3 56.1 81.5 50.6 10.3 58.7 76.7 28.8 68.7 17.2 65.2 63.1 72.1 71.7 76.2 80.0 60.9 81.7
Starcoder2-Instruct-v0.1 15B 34.2 10.0 34.3 20.0 33.3 29.2 25.3 5.2 50.7 46.7 16.9 50.7 0.0 37.7 56.9 44.3 35.0 57.1 58.6 39.1 25.0
20B+ Models
Codestral-v0.1 22B 56.1 72.9 64.2 43.3 63.8 63.1 31.3 10.3 64.0 85.0 27.1 79.1 11.7 63.8 47.7 68.9 55.0 79.4 72.9 68.8 41.7
Qwen2.5-Instruct 32B 65.8 64.3 53.7 75.0 50.7 87.7 53.0 10.3 65.3 93.3 32.2 74.6 13.3 81.2 75.4 90.2 80.0 85.7 82.9 84.4 58.3
Qwen2.5-Coder-Instruct 32B 68.2 78.6 60.9 75.0 56.1 83.1 44.6 13.8 61.3 91.7 33.9 85.1 22.4 82.6 64.6 91.8 80.0 79.4 82.9 82.8 91.7
DS-Coder-Instruct 33B 57.7 65.7 59.4 46.7 50.0 70.8 39.8 19.0 65.3 75.0 28.8 73.1 10.3 58.0 55.4 73.8 61.7 79.4 68.6 78.1 70.0
CodeLlama-Instruct 34B 28.6 70.0 23.9 18.3 26.1 15.4 18.1 10.3 40.0 18.3 25.4 46.3 3.3 24.6 24.6 49.2 11.7 60.3 38.6 14.1 25.0
Meta-Llama-3-Instruct 70B 50.1 27.1 29.9 61.7 34.8 73.8 4.8 10.3 56.0 75.0 27.1 76.1 13.3 75.4 73.8 70.5 60.0 73.0 60.0 64.1 43.3
Meta-Llama-3.1-Instruct 70B 56.6 48.6 49.3 55.0 44.9 75.4 8.4 17.2 61.3 71.7 35.6 83.6 16.7 79.7 67.7 75.4 63.3 76.2 77.1 81.2 46.7
DS-V2.5 21/236B 65.1 14.3 60.9 70.0 62.1 78.5 51.8 12.1 61.3 80.0 40.7 83.6 23.3 82.6 69.2 83.6 80.0 81.0 87.1 92.2 86.7
DS-V3 37/671B 64.9 42.9 59.7 66.7 55.1 83.1 42.2 8.6 70.7 81.7 40.7 88.1 13.3 81.2 76.9 90.0 75.0 88.9 82.9 92.2 55.0
Qwen2.5-Instruct 72B 63.6 62.9 53.7 68.3 56.5 81.5 34.9 10.3 62.7 81.7 37.3 67.2 21.7 82.6 69.2 90.2 76.7 87.3 82.9 82.8 61.7
xDebugGen (Our Method) 7B 47.5 21.4 57.8 36.7 62.1 60.0 31.3 17.2 56.0 33.3 25.4 67.2 12.1 62.3 41.5 60.7 43.3 61.9 77.1 45.3 70.0
@@
Model Size Avgall C C# CLISP CPP F# Go HTML JS Java Json Julia MD PHP Pascal Python R Ruby Rust Scala Swift
@@
Qwen2.5-Instruct 1.5B 2.0 1.4 4.5 1.7 4.3 1.5 0.0 5.2 1.3 1.7 0.0 1.5 5.0 1.4 0.0 4.9 0.0 1.6 4.3 0.0 0.0
OpenCoder-Instruct 1.5B 4.2 0.0 0.0 18.3 1.4 0.0 2.4 1.7 1.3 0.0 25.4 1.5 10.0 7.2 1.5 0.0 0.0 1.6 12.9 1.6 0.0
Qwen2.5-Instruct 3B 10.2 10.0 10.4 3.3 5.8 16.9 14.5 5.2 10.7 8.3 3.4 6.0 18.3 8.7 4.6 11.5 15.0 6.3 10.0 14.1 20.0
7B+ Models
Qwen2.5-Coder-Instruct 7B 8.4 11.4 4.7 8.3 3.0 6.2 15.7 8.6 14.7 8.3 10.2 7.5 13.8 4.3 1.5 16.4 8.3 11.1 8.6 0.0 3.3
Meta-Llama-3-Instruct 8B 3.0 2.9 4.5 3.3 2.9 0.0 3.6 0.0 8.0 6.7 1.7 0.0 0.0 4.3 0.0 9.8 1.7 0.0 7.1 0.0 1.7
Meta-Llama-3.1-Instruct 8B 5.4 7.1 10.4 3.3 11.6 0.0 7.2 1.7 5.3 3.3 1.7 6.0 3.3 8.7 3.1 9.8 1.7 6.3 4.3 1.6 10.0
Yi-Coder-Chat 9B 8.7 25.7 9.4 5.0 9.1 4.6 10.8 5.2 12.0 11.7 1.7 16.4 6.9 5.8 4.6 14.8 5.0 3.2 5.7 7.8 5.0
20B+ Models
Codestral-v0.1 22B 16.2 21.4 19.4 10.0 21.7 6.2 31.3 12.1 20.0 18.3 6.8 16.4 20.0 7.2 10.8 32.8 8.3 20.6 14.3 12.5 6.7
Qwen2.5-Instruct 32B 19.4 28.6 25.4 10.0 23.2 9.2 34.9 12.1 24 18.3 10.2 26.9 30.0 11.6 10.8 32.8 13.3 25.4 14.3 9.4 10.0
Qwen2.5-Coder-Instruct 32B 23.6 30.0 25.0 13.3 31.8 15.4 37.3 22.4 28.0 16.7 11.9 31.3 29.3 18.8 21.5 36.1 36.7 28.6 20.0 4.7 6.7
DS-Coder-Instruct 33B 12.3 14.3 15.6 0.0 15.2 6.2 15.7 12.1 22.7 8.3 6.8 11.9 27.6 10.1 10.8 24.6 6.7 17.5 11.4 4.7 1.7
Qwen2.5-Instruct 72B 17.6 28.6 16.4 13.3 18.8 7.7 25.3 19.0 26.7 15.0 8.5 20.9 28.3 13.0 6.2 26.2 21.7 22.2 12.9 7.8 10.0
DS-Coder-V2.5 21/236B 19.2 21.4 17.2 11.7 16.7 13.8 32.5 19.0 29.3 18.3 5.1 13.4 20.0 8.7 15.4 37.7 15.0 23.8 25.7 17.2 15.0
xDebugGen (Our Method) 7B 2.3 2.9 4.7 3.3 4.5 0.0 1.2 6.9 8.0 0.0 0.0 0.0 0.0 1.4 3.1 4.9 0.0 0.0 2.9 0.0 1.7
Table 5: Accuracy of different models for Bug Localization tasks on MdEval. The underlined numbers are the best scores for each language. “Avgall” represents the average scores of all code languages.
Model Size Avgall C C# CLISP CPP F# Go HTML JS Java Json Julia MD PHP Pascal Python R Ruby Rust Scala Swift
@@
Qwen2.5-Instruct 1.5B 22.8 22.9 40.3 13.3 21.7 9.2 26.5 8.6 26.7 36.7 16.9 34.3 21.7 15.9 21.5 19.7 30.0 12.7 24.3 15.6 33.3
OpenCoder-Instruct 1.5B 10.5 1.4 17.9 10.0 5.8 6.2 16.9 5.2 25.3 13.3 8.5 10.4 8.3 14.5 15.4 6.6 1.7 6.3 14.3 7.8 8.3
Qwen2.5-Instruct 3B 21.4 30.0 14.9 16.7 13.0 23.1 18.1 29.3 21.3 31.7 27.1 19.4 21.7 20.3 20.0 36.1 26.7 17.5 12.9 23.4 8.3
7B+ Models
Qwen2.5-Coder-Instruct 7B 26.8 42.9 21.9 16.7 27.3 23.1 31.3 8.6 32.0 33.3 32.2 25.4 10.3 23.2 32.3 41.0 18.3 36.5 28.6 31.2 13.3
Meta-Llama-3-Instruct 8B 7.8 8.6 6.0 3.3 5.8 1.5 10.8 12.1 10.7 11.7 18.6 6.0 18.3 4.3 3.1 3.3 1.7 6.3 11.4 10.9 1.7
Meta-Llama-3.1-Instruct 8B 7.0 7.1 9.0 10.0 5.8 3.1 12.0 17.2 0.0 6.7 27.1 4.5 13.3 2.9 1.5 3.3 3.3 0.0 8.6 7.8 0.0
Yi-Coder-Chat 9B 29.5 42.9 40.6 20.0 34.8 29.2 22.9 0.0 60.0 40.0 18.6 28.4 1.7 42.0 30.8 6.6 26.7 36.5 35.7 25.0 33.3
20B+ Models
Codestral-v0.1 22B 43.6 52.9 41.8 35.0 44.9 43.1 42.2 32.8 62.7 48.3 32.2 53.7 20.0 50.7 43.1 60.7 40.0 47.6 38.6 42.2 31.7
Qwen2.5-Instruct 32B 58.4 68.6 50.7 33.3 65.2 55.4 49.4 75.9 65.3 60.0 66.1 68.7 53.3 56.5 50.8 62.3 68.3 61.9 58.6 51.6 46.7
Qwen2.5-Coder-Instruct 32B 59.4 81.4 50.0 46.7 65.2 64.6 63.9 13.8 72.0 68.3 52.5 68.7 12.1 66.7 55.4 73.8 85.0 63.5 60.0 59.4 50.0
DS-Coder-Instruct 33B 17.8 28.6 12.5 8.3 19.7 16.9 20.5 1.7 38.7 33.3 8.5 20.9 3.4 13.0 12.3 1.6 10.0 27.0 31.4 15.6 21.7
Qwen2.5-Instruct 72B 57.4 74.3 44.8 43.3 44.9 66.2 55.4 65.5 64.0 68.3 61.0 67.2 45.0 50.7 47.7 68.9 61.7 60.3 50.0 59.4 50.0
DS-Coder-V2.5 21/236B 52.3 75.7 50.0 28.3 56.1 40.0 54.2 63.8 64.0 68.3 59.3 44.8 43.3 58.0 27.7 70.5 53.3 34.9 47.1 57.8 46.7
xDebugGen (Our Method) 7B 18.7 30.0 21.9 3.3 15.2 10.8 20.5 8.6 30.7 36.7 28.8 23.9 8.6 18.8 15.4 13.1 5.0 33.3 20.0 15.6 6.7
@@ 3.2 Main Results
Table 3 presents the Pass@1 results of different models on MdEval for the multilingual automated program repair task (given question with buggy code, request the model to fix buggy code). The results indicate a marked disparity between closed-source state-of-the-art models and the majority of open-source models across nearly all programming languages. Notably, GPT-4o, Claude-3.5-sonnet, and Qwen2.5-Coder-Instruct excel in this task and demonstrate significant performance advantages over other models. Furthermore, our baseline model xDebugCoder, is fine-tuned using only 16K bug-related data MdEval-Instruct. Despite the limited size of this dataset, the model demonstrated competitive performance compared to others of similar scale, highlighting the effectiveness of MdEval-Instruct in enhancing the debugging capabilities of models.
Bug Localization
Table 5 illustrates the accuracy of different models on the multilingual bug localization task. It is evident that closed-source models outperform open-source models by a significant margin, demonstrating the superior bug localization capabilities of closed-source models. Specifically, open-source models with smaller parameter sizes like OpenCoder-1.5B-Instrcut, due to their poor instruction-following capabilities, are unable to output the correct format as required, resulting in lower accuracy in localization. Besides, it is observed that for the same model, the bug localization accuracy is lower than its pass@1 scores in automated program repair task. We hypothesize that this discrepancy arises because the bug localization task requires a strong understanding of location information, which happens to be a weakness of large language models. Therefore, improving LLMs’ ability to understand location information is a critical issue that needs to be addressed.
@@
In the bug identification task, the goal is to identify the error type in a given code snippet, where the LLMs analyze source code for defects and choose the correct bug type from the pre-defined 47 bug types. Table 4 lists the all results of the bug identification. Notably, the closed-source LLMs, such as GPT-4o and Claude series, have the dominant advantages, outperforming the open-source LLMs by nearly +10 points. Bug identification with 47 bug types poses a daunting challenge to the LLMs, requiring alignment capability of LLMs between the given code snippet and its corresponding bug type. As a result, some open-source models with smaller parameter sizes perform poorly in this task
@@ 4 Further Analysis
In Figure 6, The performance of models on the automated program repair task varies across different error types, highlighting the strengths and weaknesses of these models in addressing specific challenges. Consistently, the models demonstrate robust capabilities in repairing syntax errors, reference errors, and logic errors. These error types tend to be more straightforward and well-defined, allowing the models to leverage their knowledge effectively to identify and correct issues with high accuracy. In contrast, the models exhibit their worst performance when dealing with language-specific errors. Language-specific errors can arise from unique syntax rules, idiomatic expressions, or even cultural programming practices that are not universally applicable. As a result, addressing these types of errors presents a significant challenge and underscores the need for further improvements in model training.
@@ 4.2 Another Two Automated Program Repair Settings
4.3 Anslysis of Code Review task
Besides automated program repair tasks, code review tasks also play a crucial role in software development. To analyze the performance of different models on code review tasks, we conducted experiments based on MdEval. For the code review task, we present two versions of code to LLMs: the correct code $b^{L_{k}}$ and the buggy code $c^{L_{k}}$ with only a few minor differences between them. The correct code and buggy code are listed in a random order to feed into LLM for distinguishing the buggy code.
Figure 8 displays the accuracy for code review tasks. The results show that closed-source models still significantly outperform open-source models in the code review task. The closed-source models demonstrate a strong ability to understand complex code logic, achieving an accuracy rate of approximately 90%. In contrast, the smaller open-source model exhibits significant challenges, with an accuracy rate of around 50%. This disparity underscores the limitations of the current open-source model in effectively interpreting intricate coding patterns.
Figure 8: Accuracy of different models for Code Review tasks on MdEval.
4.4 Analysis on the Effect of Bug Location for Automated Program Repair
In previous studies, bug localization has been regarded as the first step in program repair, playing a critical role. To verify whether the bug location information can also have a positive impact when using large language models for automated program repair, we designed and conducted a series of comparative experiments, as shown in  Figure 9. We test two scenarios: providing the bug location information and not providing it and task the model with repairing buggy code in both cases. The results indicate that providing the bug location information significantly improves Pass@1 scores of automated program repair. However, our prior experimental results reveal that for LLMs, the difficulty of the bug localization task is notably higher than that of the automated program repair task. Therefore, improving the bug localization capabilities of the model is essential for enhancing its overall automated program repair performance.
Figure 9: Comparison of the Pass@1 (%) scores with only the buggy code provided versus when additional bug location information is supplied.